OpenStudy (xapproachesinfinity):

here is a fun one: if \[x\sin \pi x=\int_{0}^{x^2}f(t)dt\], where f is continuous find f(4)

Nnesha (nnesha):

hmmm very funny question seem like fun fun!!!!!!!!!!!!!1

OpenStudy (xapproachesinfinity):

isn't it fun ? @Nnesha haha

OpenStudy (xapproachesinfinity):

@freckles would like this one but offline now darn ^^

Nnesha (nnesha):

nope i don't think so

OpenStudy (anonymous):

Not an Infinite Series

OpenStudy (anonymous):

Not an Infinite Series

OpenStudy (xapproachesinfinity):

okay let me see my attempt: \[\frac{d}{dx}(x\sin \pi x)=\frac{d}{dx}\int_{0}^{x^2}f(t)dt\] \[\sin \pi x+\pi x\cos \pi x=2xf(x^2)\] \[f(x^2)=\frac{\sin \pi x+\pi x\cos \pi x}{2x}\] \[f(2^2)=\frac{\sin2\pi +2\pi \cos 2\pi}{4}\] \[f(4)=\frac{2\pi}{4}=\frac{\pi}{2}\] \[\square\]

OpenStudy (xapproachesinfinity):

@Nnesha see quite fun one :)

OpenStudy (xapproachesinfinity):

@perl check if this is correct

OpenStudy (perl):

yes i agree

OpenStudy (perl):

nice work :)

OpenStudy (xapproachesinfinity):

thanks :)

OpenStudy (xapproachesinfinity):

this was in problem plus section LOL i'm falling in love with it heheh